Consider the hypotheses shown below. Given that x overbar= 48 ​,sigma=13 ​, n=33​, alpha=0.10 ​, complete parts a and b.

Upper H 0​: mu less than or=45

Upper H 1​: mu greater than 45

a. What conclusion should be​ drawn?

b. Determine the​ p-value for this test.

a. The​ z-test statistic is .

​(Round to two decimal places as​ needed.)
Answer #1

Given that, sample size ( n ) = 33

sample mean (ar x) = 48

population standard deviation (sigma) = 13

significance level (alpha) = 0.10

The null and the alternative hypotheses are,

Ho : μ < 45

1 u > 45

Test statistic is,

2. _ μ 48-45 Vn y 33

Test statistic is 1.33

p-value = P(Z > 1.33) = 1 - P(Z < 1.33) = 1 - 0.9082= 0.0918

p-value = 0.0918

Here, p-value = 0.0918 < α = 0.10

So, we reject the null hypothesis ( H0).